OpenStudy (anonymous):

:) 0=x^2+6x-16 what is the descrimanit?

OpenStudy (anonymous):

x^2-6x-16=0 (x+2)(x-8) = 0 x+2 = 0 or x-8 = 0 x = -2 or x = 8 I hope this helps!